# mode=buildnew

### v2

@@ 3.2.2 Multi-objective optimization via shrinking the reference signal in constrained optimization
Remark 3.4.
In the case of $R(\cdot)\in\mathbb{R}^{d>1}$, when all components of $\mu$ increases, Equation 32 and Equation 33 imply at least one component of $R(\cdot)$ will decrease. We defer to Remark 3.6 and Section 3.2.3 for the discussion on when some components of $\mu$ increases while other components decreases.
Remark 3.5.
Definition 3.4 and Corollary 3.1 gives a condition of decreasing $R(\cdot)$ at the expense of increasing $\ell(\cdot)$ (increase bounded though) at iteration $k$. When the gradient component from $\ell(\cdot)$ and $R(\cdot)$ agrees with each other at another iteration, we could expect both objectives to decrease as in Definition 3.3. It can also be the case at a particular iteration, $\ell(\cdot)$ decreases at the expense of $R(\cdot)$ increases.
What is left unclear, however, is after the accumulative effect of several iterations, if $\exists m$ to ensure Equation 30. To answer this, first, Corollary 3.2 gives a bound on the accumulative change of $\ell(\cdot)$ with $m=B$.
Corollary 3.2.
Suppose that for all $k=0,\dots,B-1$
and $\mu^{(k+1)}>\mu^{(k)}$. Then
$\ell((\theta^{(B)})\leq\ell(\theta^{(0)})+S_{>}+S_{<},$ (44)
where
$S_{<}$ $=\sum_{k\in\mathcal{K}_{<}}\mu^{(k)}(R(\theta^{(k)})-R(\theta^{(k+1)})),$
[…]
$\mathcal{K}_{<}$ =$=\{0,\dots,B-1\}\setminus\mathcal{K}_{>}.$
@@
For $k\in\mathcal{K}_{>}$ (Equation 33 holds), from Corollary 3.1 we have that
@@
$\ell((\theta^{(B)})-\ell(\theta^{(0)})$
$=\sum_{k=0}^{B-1}[\ell((\theta^{(k+1)})-\ell(\theta^{(k)})]$
$=\Big{(}\sum_{k\in\mathcal{K}_{>}}+\sum_{k\in\mathcal{K}_{<}}\Big{)}[\ell((\theta^{(k+1)})-\ell(\theta^{(k)})]$
yields the conclusion. ∎
Based on Corollary 3.2, we have the following conjecture to ensure Equation 30
Conjecture 3.1.
Further decompose $S_{<}$ from Corollary 3.2 into $S_{<}^{-}$ and $S_{<}^{+}$ via decomposing $\mathcal{K}_{<}$ in Corollary 3.2 into $\mathcal{K}_{<}^{-}$ and $\mathcal{K}_{<}^{+}$. Here $\mathcal{K}_{<}^{-}$ corresponds to the situations of $\ell(\cdot)$ descent with $S_{<}^{-}<0$. And $\mathcal{K}_{<}^{+}$ corresponds to the $\ell(\cdot)$ ascent cases but upperbounded by $S_{<}^{+}=\sum_{k\in\mathcal{K}_{<}^{+}}\mu^{(k)}(R(\theta^{(k)})-R(\theta^{(k+1)}))$. Then $\exists$ sequence $\{\mu^{(k)}\}$ s.t.
$\ell((\theta^{(B)})\leq\ell(\theta^{(0)})+S_{>}+S_{<}^{+}+S_{<}^{-}<0$ (45)
Remark 3.6.
In the case of $d>1$, as long as a loss component of $R(\cdot)^{(k)}$, say $R_{1}^{(k)}(\cdot)$ is not bounded by its corresponding setpoint component $b_{1}^{(k)}(\cdot)$, we can still increase the value of the $\mu$ component $\mu^{(k)}_{1}$ to give more weight to the corresponding loss component, until the loss component in question decreases below the setpoint. However, an increased $\mu$ component $\mu^{(k)}_{1}$ corresponds to more weight on the gradient component corresponding to $R_{1}^{(k)}(\cdot)$, which makes it harder for other components of $R(\cdot)$ and $\ell(\cdot)$ to decrease. Therefore, instead of always increasing each component of $\mu$, if one component of $R$ has overshoot over the setpoint (i.e. constraint switched from being not satisfied to satisfied with respect to a particular loss component), we decrease the $\mu$ value to give the other components of $\mu$ and $R$ more feasible space to adjust themselves.
Remark 3.7.
We could extend the conjecture above to $\exists\mu$ sequence such that after $B$ iterations, Pareto descent in Definition 3.3 could be achieved (equivalently, setpoint should shrink see Remark 3.3). We did observe such a Pareto descent behavior, defined in Definition 3.3 in our experiment for the $d>1$ case of Equation 4, e.g. see Figure 5 and Figure 6.
3.2.3 Output feedback PI-like controller
How exactly should the $\mu$ sequence discussed above look like? With the adaptive law from Equation 23 for setpoint $b$, we design the following controller for $\mu$:
$e^{(k)}$ $=R(\cdot)-b^{(k)}$ (46)
$\delta_{I}^{(k+1)}$ $=(1-\xi_{d})\delta_{I}^{(k)}+\xi_{d}e^{(k)}$ (47)
$K_{I}$ $>0$ (49)
$K_{I}$ $\in\mathbb{R}^{d}$ (50)
$k$ $=0,1,\cdots,B$ (51)
In Equation 46, we calculate how far the current output is away from the setpoint $b$, which gets passed through a moving average in Equation 47 with $\xi_{d}$ being the coefficient of moving average in Equation 47 (Rezende and Viola, 2018). $\mu$ is the output of the controller in Equation 48, $K_{I}$ is the control gain for PI (proportional-integration)  Johnson and Moradi (2005) like control, $K_{I}\delta_{I}^{(k+1)}$ is component wise multiplication, $\exp$ and $\max$ are computed component wise, $v_{sat}$ is the exponential shoulder saturation, $\mu_{clip}$ is the upper bound for $\mu$. Note that $\mu_{clip}$ determines the dynamic range of $\mu$ while $v_{sat}$ determines an upper bound about how fast $\mu$ changes.
Remark 3.8.
The whole process is summarized in Algorithm 1 and Figure 1.
1:procedure M-Hof-Opt($\mu^{(0)},\theta^{(0)},\rho,\cdot$)
2:     Initialize $\mu^{(0)}$, calculate $b^{(0)}=\rho R^{(0)}$
3:     Compute $K_{I}$ based on Remark 3.8
4:     while budget $B$ not reached do
5:         Update $\mu$ from controller in Equation 48
6:         Update $\theta$ via running Equation 14 with the updated $\mu$ value
7:         Adapt setpoint according to Equation 23.
8:     end while
9:     return $\theta^{(B)}$triangleright
10:end procedure
3.2.4 Multi-objective setpoint based model selection
4 Related work and Discussion
@@
Multipliers warmup (Sønderby et al., 2016; Ilse et al., 2020) or phase-in (Sicilia et al., 2023) requires ultimate multiplier values for each loss term, thus necessitates hyperparameter (ultimate multiplier values) optimization. Other works, such as (Rezende and Viola, 2018; Klushyn et al., 2019), has introduced constraint optimization by representing the weights of loss terms using Lagrange multipliers, designed to prevent over-regularisation of a single term within the loss function, which is extended to multiple loss terms in (Chen et al., 2022a). In addition, Shao et al. (2020) (only deals with 2 loss terms),Chen et al. (2022a) proposed similar multiplier adaptation schemes as in Section 3.2.3 where the multiplier changes with a rate proportional to how the corresponding loss component distances from a constraint bound. However, their methods relied on a fixed constraint bound  for each loss term and fixed proportionate gain for each loss term (defined to be $K_{I}$ in Equation 48 in our work), necessitating a hyperparameter search, which is challenging when the number of terms in the loss increases. Moreover, for some regularizers, such as the ones based on KL divergence, it is challenging to estimate the constraint bound range. Specifically in the field of domain generalization, it is not obvious how to choose the constraint bound value and other hyperparameters due to the lack of observations from the target domain, where the validation set from the training domain can easily reach a saturated value due to the high expressive power of modern neural networks.
@@
In comparison, we deal with many loss terms with multidimensional multipliers and introduced Equation 23 to progressively adapt the constraint bound and Remark 3.8 to choose the multidimensional controller gain value, thus avoid exhaustive hyperparameter search. In addition to the control theory formulation for multiplier adjustment, we also provided a Probabilistic Graphical Model interpretation in Section 3.1 and a landscape scheduling interpretation in Figure 4. Furthermore, we discussed our method under the multi-objective optimization scheme compared to single-objective consideration from other works.
@@
Earlier multi-objective deep learning (Sener and Koltun, 2018; Lin et al., 2019) suffers from increased computational cost with the number of losses, explicit computation of the gradient with respect to each loss term and gradient norms, making it practically infeasible Mahapatra and Rajan (2021) for large neural networks and large datasets. For instance, Pareto Invariance Risk Minimization (PAIR) (Chen et al., 2022b) studies domain generalization with the use of the multi-objective optimization technique Exact Pareto Optimal Search (EPOS) (Mahapatra and Rajan, 2021), which requires explicit full gradient information for each loss term and computation of the $C$ matrix with complexity $m^{2}n$ (Mahapatra and Rajan, 2021). With $m$ being the number of loss terms (equivalent to $d+1$ in our paper), and $n$ the dimension of gradient, which is super high in modern neural networks (e.g. 200 million for ResNet50), thus results in heavy memory requirements and excessive computation burden, and hinders the deployment to large-scale networks Chen et al. (2022b). In comparison, our multi-objective optimization method (M-HOF-Opt) need only $d$ scalar multiplications, thus the extra computation is negligible and can be used in large-scale neural networks. EPOS (Mahapatra and Rajan, 2021) also requires the user to provide a preference vector which can be difficult to set, say with 6 objectives for our case. In addition, Chen et al. (2022b) reported hyperparameter tuning for the preference vector and step length is needed for the method to work properly. In comparison, our method, M-HOF-Opt is preference vector free. It ensures the Pareto descent due to the update rule of our setpoint. There is no requirement to search for step length.
@@
5 Experiments
Figure 5: Our method drives the different components of $R(\cdot)$ loss terms down at different scales and rates towards the setpoint, which further promotes the setpoint shrinkage. In the Top row, we show the multiplier dynamic as controller output signal in Equation 48 with respect to training epochs. In the Bottom row, we present the tracking behavior of the corresponding regularization loss $R(\cdot)$ in Equation 4 with respect to setpoint $b$ defined in Equation 23. This figure is from training DIVA (Ilse et al., 2020) with our multi-objective hierarchical feedback optimization scheme on the PACS dataset (Li et al., 2017). Left column: For the regularization loss of negative KL-divergence corresponding to multiplier $\beta_{y}$ in  Equation 8 with $R_{\beta_{y}}=-KL(q_{\phi_{y}}(z_{y}|x)||p_{\theta_{y}}(z_{y}|y))$ (blue curve), we see that with the setpoint (red curve) initially set to $\rho=0.99$ in Equation 21 of the initial $R_{\beta_{y}}$ value, in the first few epochs $R_{\beta_{y}}$ overshoots to an even worse value, but our controller pulls it back towards the setpoint in later epochs via increase in $\beta_{y}$ (top figure in the same column) thus adding more weight on $R_{\beta_{y}}(\cdot)$ in the penalized loss. When $R_{\beta_{y}}$ is controlled to be bounded by its setpoint, the corresponding multiplier value $\beta_{y}$ also decays to give space for other objectives to decrease (top figure in the same column). With coordinated $\mu$ dynamic from each component of the controller output, the setpoint shrinks when all $R(\cdot)$ components Pareto dominates the setpoint. After the setpoint shrinks several times, feasibility in Equation 26 becomes harder to maintain and our controller still keeps $R_{\beta_{y}}$ oscillating around the setpoint. Our model selection criteria in Section 3.2.4 selects the model at the last occurrence of the setpoint shrinkage, thus the oscillating behavior at the later epochs can be ignored. Middle column: For the domain classification loss $R_{\gamma_{d}}$ corresponding to multiplier $\gamma_{d}$ in  Equation 10 (blue curve), we observe a similar behavior, but the setpoint shrinkage is more visible. Right column: Although the output component for the loss $R_{\mu_{recon}}$ in Equation 6 oscillates with relatively large amplitude after the last setpoint shrinkage, our model selection criteria at Section 3.2.4 will exclude this oscillating behavior.
Figure 6: Our training scheme adapts the multi-dimensional multiplier $\mu$, steering the optimization process towards a configuration that balances the trade-off between minimizing output $\ell(\cdot)=\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]$ in Equation 5 and output $R_{\gamma_{d}}(\cdot)=\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ in  Equation 10. In this plot, the solid large point corresponds to the initial output. We use gradually changing colors to indicate the training iterations (epochs), indicated by the color bar. For improved visualization, we exclusively utilized the initial 120 epochs and plotted data points solely for every 10 epochs. This figure corresponds to Figure 5.
5.1 Experimental Setting
Figure 7: Our automatic multiplier adjustment scheme ensures robust out-of-domain test accuracy when training DIVA (Ilse et al., 2020) with the modified ResNet50 from (Gulrajani and Lopez-Paz, 2020) on PACS dataset (testing on domain sketch and training on domain photo, art-painting and cartoon). Each panel (i.e., subplot) corresponds to a training scheme. Each box plot inside a panel corresponds to a specific hyperparameter (e.g. multiplier value or controller hyperparameters) combination. Inside each box plot, we repeat the experiment with different random seeds corresponding to the dots scatter. In the first panel, we warm up only $\beta_{y}$, $\beta_{d}$ in Equations 9 and 8 to their ultimate value $1$ while sampling fixed multiplier $\gamma_{d}$ and $\gamma_{y}$ while fixing $\mu_{recon}=1.0$. In the second panel, additionally, we also warm up $\mu_{recon}$ in Equation 6 to $1.0$ and warm up $\gamma_{d}$ in Equation 10 to the sampled value, and keep $\gamma_{y}$ fixed to be the sampled value. In the third panel, we use fixed constant multipliers without warm-up while keep $\beta=1$ and $\mu_{recon}=1$ and only sample $\gamma_{d},\gamma_{y}$ combinations. The last two panels correspond to results utilizing our multi-objective hierarchical feedback optimization training method, and we let $\mu$ include all multipliers except $\gamma_{y}=1$ in Equation 5. For simplicity, when we write $\mu^{(0)}=1.0$ where $\mu^{(0)}\in\mathbb{R}_{+}^{d}$, we mean each component of $\mu^{(0)}$ equals $1.0$.
@@ 5.1 Experimental Setting
This section presents our experiments training the domain generalization model DIVA (Ilse et al., 2020) with $6$ loss terms defined in Equation 4 on the widely used domain generalization benchmark dataset PACS (Li et al., 2017) with different training schemes. Due to the excessive memory requirements and heavy computation of earlier multi-objective deep learning methods discussed in Section 4, which hinders their application, we can only compare our training scheme with other multiplier scheduling techniques. Given that the sketch domain within the PACS dataset is considered the most inherently challenging, it was chosen as the sole leave-one-out domain to test the performance of domain generalization. We run experiments to answer the following questions:
  • Is our multiplier controller capable of adjusting multidimensional multipliers automatically to drive each component of $R(\cdot)$ at different scales to the setpoint? Does the setpoint shrink, which implies multi-objective descent? We answer these questions with Figure 5 and Figure 6.
  • Will wrong multiplier choices for warmup and fixed multiplier training schemes result in catastrophic effects? Although our method in contrast is multiplier-free, it does need multiplier initial condition $\mu^{(0)}$ and controller hyperparameters. Does varying those have detrimental effects on the predicative performance as well? We answer this question with the benchmark setting in Section 5.1.1 and the results presented in Figure 7.
5.1.1 Benchmark setting

### v3

@@ Appendix A Mathematical proofs
See 3.3
@@
For $k\in\mathcal{K}_{>}$ (Equation 21 holds), from Proposition 3.2 we have that
@@
$\ell(\theta^{(B)})-\ell(\theta^{(0)})$
$=\sum_{k=0}^{B-1}[\ell(\theta^{(k+1)})-\ell(\theta^{(k)})]$
$=\Big{(}\sum_{k\in\mathcal{K}_{>}}+\sum_{k\in\mathcal{K}_{<}}\Big{)}[\ell(\theta^{(k+1)})-\ell(\theta^{(k)})]$
yields the conclusion.
Appendix B Related work and discussion
In this section, we discuss the advantages of our method compared to other multiplier scheduling and multi-objective optimization methods, especially in time and space (memory) complexity. Compared to hyperparameter tuning including Bayesian optimization as discussed in Remark 3.2, since in general the next multiplier has to be generated after one training round, in the extreme worst case, the number of searches reduces to grid search of $B_{d}^{d}$ number of combinations (where $B_{d}$ is the number of grid points for each multiplier component, $d$ is the dimension of the multiplier). Our method however requires only a single run.
B.1 Multiplier adaptation in deep learning
@@
Multipliers warmup (Sønderby et al., 2016; Ilse et al., 2020) or phase-in (Sicilia et al., 2023) requires ultimate multiplier values for each loss term, thus necessitates hyperparameter (ultimate multiplier values) optimization, see also analysis in Remark C.1. Other works, such as (Rezende and Viola, 2018; Klushyn et al., 2019), has introduced constraint optimization by representing the weights of loss terms using Lagrange multipliers, designed to prevent over-regularisation of a single term within the loss function, which is extended to multiple loss terms in (Chen et al., 2022a). In addition, similar multiplier adaptation schemes were proposed by (Shao et al., 2020), who only dealt with two loss terms, and by (Chen et al., 2022a), where the multiplier changes at a rate proportional to the distance of the corresponding loss component from a setpoint, as discussed in Section 3.2.5. However, their methods relied on a fixed constraint bound for each loss term and fixed proportionate gain for each loss term (defined to be $K_{I}$ in Equation 31 in our work), necessitating a hyperparameter search, which is challenging when the number of terms in the loss increases. Moreover, for some regularizers, such as the ones based on KL divergence, it is challenging to estimate the constraint bound range. Specifically in the field of domain generalization, it is not obvious how to choose the constraint bound value and other hyperparameters due to the lack of observations from the target domain, where the validation set from the training domain can easily reach a saturated value due to the high expressive power of modern neural networks.
@@
In comparison, we deal with many loss terms with multidimensional multipliers and introduced Equation 14 to progressively adapt the constraint bound and Remark 3.13 to choose the multidimensional controller gain value, thus avoid exhaustive hyperparameter search. In addition to the control theory formulation for multiplier adjustment, we also provided a probabilistic graphical model interpretation in Section 3.1 and a landscape scheduling interpretation in Figure 4. Furthermore, we discussed our method under the multi-objective optimization scheme compared to single-objective consideration from other works. Different from the single layer controller in (Chen et al., 2022a), we proposed a hierarchical controller in Section 3.2.5 to break the multi-objective optimization problem into a series of constraint optimization sub-problems, each configured with a self-adaptive multi-objective setpoint updated via Pareto dominance.
B.2 Multi-objective optimization for large-scale neural networks
@@
Earlier multi-objective deep learning (Sener and Koltun, 2018; Lin et al., 2019) suffers from increased computational cost with the number of losses, explicit computation of the gradient with respect to each loss term and gradient norms, making it practically infeasible Mahapatra and Rajan (2021) for large neural networks and large datasets. For instance, Pareto Invariance Risk Minimization (PAIR) (Chen et al., 2022b) studies domain generalization with the use of the multi-objective optimization technique Exact Pareto Optimal Search (EPOS) (Mahapatra and Rajan, 2021), which requires explicit full gradient information for each loss term and computation of the $C$ matrix with complexity $m^{2}n$ (Mahapatra and Rajan, 2021). With $m$ being the number of loss terms (equivalent to $d+1$ in our paper), and $n$ the dimension of gradient, which is super high in modern neural networks (e.g. 200 million for ResNet50), thus results in heavy memory requirements and excessive computation burden, and hinders the deployment to large-scale networks (Chen et al., 2022b). The Dirichlet sampling of preference vectors was introduced, and the aggregated optimal was computed under the sampled preference vectors (Ruchte and Grabocka, 2021). However, their sampling of preference vectors introduced data replications. In addition, their method needs to augment the preference vector with the original input, which added complexity of applying their method to different data modes. For instance, for image input, they need first to transform the preference vector to image mode via transpose convolution. Further, their method needs an extra hyperparameter to weight the additional penalty loss to force the solution to obey the preference vectors.
In comparison, our multi-objective optimization method (M-HOF-Opt) need only $d$ scalar multiplications, thus the extra computation is negligible and can be used in large-scale neural networks. EPOS Mahapatra and Rajan (2021) also requires the user to provide a preference vector which can be difficult to set, say with 6 objectives for our case. In addition, Chen et al. (2022b) reported hyperparameter tuning for the preference vector and step length is needed for the method to work properly. In comparison, our method, M-HOF-Opt is preference vector free. It ensures the Pareto descent due to the update rule of our setpoint. There is no requirement to search for step length.
@@
Appendix C Experimental details
C.1 Experimental Setting
@@
This section presents our experiments training the domain generalization model DIVA (Ilse et al., 2020) with $6$ loss terms defined in Equation 40 on the widely used domain generalization benchmark dataset PACS (Li et al., 2017) with different training schemes. Due to the excessive memory requirements and heavy computation of earlier multi-objective deep learning methods discussed in Appendix B, which hinders their application, we compare our training scheme with other multiplier scheduling techniques. Given that the sketch domain within the PACS dataset is considered the most inherently challenging, it was chosen as the sole leave-one-out domain to test the performance of domain generalization.
To further show the power of M-HOF-Opt in automatically balancing different loss terms, we combined two domain generalization algorithms IRM (Arjovsky et al., 2019) and DIAL (Levi et al., 2021).
In our experiment, we used the Facebook Research version of ResNet from DomainBed [^2] https://github.com/facebookresearch/DomainBed/blob/main/domainbed/networks.py (Gulrajani and Lopez-Paz, 2020) with a learning rate of 5e-5 and batch size of 32.
We first introduce domain generalization and detail the loss formulation of the above mentioned methods and their loss combination, then we present the benchmark setting in Section C.3Section C.4 and the corresponding experimental results.
C.2 Domain generalization structral risk
C.2.1 Domain Invariant Variational Autoencoding (DIVA)
$L(\theta,\mu,x,y,d)=\ell(\cdot)+\mu^{T}R(\cdot)$ (40)
$=\gamma_{y}\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]+$ (41)
$\mu_{recon}\mathbb{E}_{q_{\phi_{d}}(z_{d}|x),q(z_{x}|x),q(z_{y}|x)}\log p_{\theta_{r}}(x|z_{d},z_{x},z_{y})$ (42)
$\quad+[-\beta_{x}KL(q_{\phi_{x}}(z_{x}|x)||p_{\theta_{x}}(z_{x}))]$ (43)
$\quad+[-\beta_{y}KL(q_{\phi_{y}}(z_{y}|x)||p_{\theta_{y}}(z_{y}|y))]$ (44)
$\quad+[-\beta_{d}KL(q_{\phi_{d}}(z_{d}|x)||p_{\theta_{d}}(z_{d}|d))]$ (45)
$\quad+\gamma_{d}\mathbb{E}_{q_{\phi_{d}}(z_{d}|x)}[\log q_{\omega_{d}}(d|z_{d})]$ (46)
In (Ilse et al., 2020), $\mu_{recon}=1.0$ in Equation 42. Additionally, $\gamma_{y}$ and $\gamma_{d}$ are maintained as constants. $\gamma_{y}$ is corresponding to $\ell(\cdot)=\gamma_{y}\mathbb{E}_{q_{\phi_{y}}(z_{y}|x)}[\log q_{\omega_{y}}(y|z_{y})]$ in Equation 41, and $\gamma_{d}$ in Equation 46 is associated with one component of $R(\cdot)$. However, the combined choice of $\gamma_{y}$ and $\gamma_{d}$ significantly influences the generalization performance, as shown by our experimental findings in Figure 7.
Remark C.1 (feedforward scheduler).
C.2.2 Invariant Risk Minimization (IRM) and Domain Invariant Adversarial Learning (DIAL)
Let $\Phi$ be the feature extraction neural network, $w$ be the task network (e.g. classifier), let $d$ index available training domains. Invariant Risk Minimization (IRM) uses the following optimization:
$\min_{\Phi,w}\sum_{d}\ell^{(d)}(w\circ\Phi(X))+$
$\mu\sum_{d}\|\nabla_{w|w=1.0}\ell^{(d)}(w\circ\Phi(X))\|^{2}$ (47)
while Domain Invariant Adversarial Learning proposed using adversarial perturbed image ($X_{adv}$ in comparison to original data $X$) as new training samples to form adversarial loss. If we combine these two, we essentially get the following loss:
$\ell+\mu^{T}R$
$=$ $\min_{\Phi,w}\sum_{d}\ell^{(d)}(w\circ\Phi(X))+$
$\mu_{1}\sum_{d}\|\nabla_{w|w=1.0}\ell^{(d)}(w\circ\Phi(X))\|^{2}+$
$\mu_{2}\ell(w\circ\Phi(X_{adv}))$ (48)
C.3 Benchmark settings for DIVA